Rewrite the equation –8x – 5y = –45 as a function of x.
Answer by edjones(8007)

–8x – 5y = –45
-8x+8x-5y=8x-45
-5y=8x-45
-5y/-5=8x/-5-45/-5
y=-8x/5 +9
Ed